### v1

# Goldilocks Isolation: High Performance VMs with Edera

Abstract
Organizations run applications on cloud infrastructure shared between multiple users and organizations. Popular tooling for this shared infrastructure, including Docker and Kubernetes, supports such multi-tenancy through the use of operating system virtualization. With operating system virtualization (known as containerization), multiple applications share the same kernel, reducing the runtime overhead. However, this shared kernel presents a large attack surface and has led to a proliferation of container escape attacks in which a kernel exploit lets an attacker escape the isolation of operating system virtualization to access other applications or the operating system itself. To address this, some systems have proposed a return to hypervisor virtualization for stronger isolation between applications. However, no existing system has achieved both the isolation of hypervisor virtualization and the performance and usability of operating system virtualization.
We present Edera, an optimized type 1 hypervisor that uses paravirtualization to improve the runtime of hypervisor virtualization. We illustrate Edera’s usability and performance through two use cases. First, we create a container runtime compatible with Kubernetes that runs on the Edera hypervisor. This implementation can be used as a drop-in replacement for the Kubernetes runtime and is compatible with all the tooling in the Kubernetes ecosystem. Second, we use Edera to provide driver isolation for hardware drivers, including those for networking, storage, and GPUs. This use of isolation protects the hypervisor and other applications from driver vulnerabilities. We find that Edera has runtime comparable to Docker with .9% slower cpu speeds, an average of 3% faster system call performance, and memory performance 0-7% faster. It achieves this with a 648 millisecond increase in startup time from Docker’s 177.4 milliseconds.
1 Introduction
Cloud computing allows applications to share infrastructure, reducing the cost of deployment. By offloading the running of applications, organizations can use hardware more efficiently, scale resource usage up and down with traffic, and provision their software quickly. In the early days of cloud computing, applications were run on virtual machines [52], which gave each application its own operating system (OS) kernel. Developers could run applications in this virtual machine, with a hypervisor managing the hardware sharing between many virtual machines. Now, it is more common to run multiple applications on the same kernel using operating system virtualization (OS virtualization) to separate the runtimes of each application [36, 51, 50]. This transition is due to both the efficiency of OS virtualization and the improved ease of use from a large ecosystem of tools available to use OS virtualization to run containers.
Containers are an abstraction over applications used by orchestration frameworks like Docker and Kubernetes to facilitate the use of OS virtualization. In this abstraction, the OS kernel runs a container engine that manages the creation and operation of multiple isolated runtimes, known as containers. This is a powerful abstraction that has been used to add efficiency, monitoring, and security to OS virtualization. By 2027, Gartner predicts that 90% of organizations will be running containerized applications in production [12]. Many of these organizations use OS virtualization in multi-tenant environments, with containers from different users running on the same host OS. A company may set up a collection of containers on a host OS to manage multiple copies of their service, with each customer assigned to one of these copies. Although resource sharing enabled by OS virtualization has improved application runtime and resource usage, it comes at the cost of strong isolation between applications.
OS virtualization uses kernel isolation techniques to isolate applications. These techniques, including namespaces [34], cgroups [33], seccomp [35], and capabilities [32], limit an application’s access to kernel resources such as system calls, privileges, and hardware to allow the application to run independently of other containers. However, the nature of the shared kernel leaves a large attack surface for container escapes in which an attacker running an application in a container is able to gain access to the kernel or other containers, “escaping” the isolation of the container runtime. If a cloud service provider gives a container to each customer, a container escape would allow an attacker to access or manipulate other customers’ sensitive user data. These attacks are frequent [1, 24, 43, 15, 16], with five novel container escapes found in 2022 [9, 7, 6, 10, 8].
In order to prevent container escapes, virtualized applications need stronger isolation, without sacrificing the runtime and ease-of-use present in modern virtualization systems. We find that container escape vulnerabilities include a kernel exploit that an attacker uses to circumvent kernel isolation techniques. These attacks can be prevented by moving the shared kernel out of the trusted computing base (TCB). Hypervisor virtualization, which was used for cloud computing before the creation of OS virtualization technologies in the 2000s, provides a kernel for each application, providing stronger application isolation. A hypervisor manages the virtual machines and provides access to hardware resources. By eliminating the shared kernel, an attacker who compromises the kernel will not gain access to the host machine or other applications without also compromising the host machine’s hypervisor. However, any solution that provides strong isolation will not be adopted unless it can match the runtime and ease-of-use of OS virtualization.
Prior work in improving container isolation has made progress toward achieving this combination of strong isolation and ease-of-use. Unikernel approaches [24, 43, 19, 47, 37] combine the application and kernel into a single light-weight workload. However this approach generally requires specifically building applications for use in a unikernel, as well as re-building on every iteration of the application, hindering development speed. Other virtual machine approaches either require specialized hardware [29, 1, 15, 48, 25, 26], such as CPUs that support virtualization extensions, or have high performance costs [27, 1] due to system call indirection. While these approaches have been used for security-critical applications, these limitations have prevented their widespread adoption.
In this work, we observe that paravirtualization [56, 4] can improve the performance of hypervisor virtualization. Paravirtualization reduces the overhead of hypervisors by requiring that guest OSs replace some system calls with hypercalls to the hypervisor, avoiding costly emulation of I/O components. Paravirtualization has been optimized for cloud use cases [23, 5]. We further observe that modern programming language features can reduce the attack surface of a hypervisor by reducing the occurrence of memory bugs in code. Together, these observations allow us to build an optimized hypervisor for running isolated applications.
Based on these observations, we present Edera, a paravirtualized type-1 hypervisor that provides strong isolation between applications, eliminating container escapes. Edera introduces zones that encapsulate each isolated area of the system. Edera uses a microkernel hypervisor running in a hardened root zone to control memory and CPU usage, then runs applications in zones that act as virtual machines, each containing a full Linux kernel. The microkernel, written in MISRA C, presents a small attack surface for applications running within the zones. The runtime, CPU, and memory usage of Edera are similar to that of popular OS virtualized systems, with less than a second of additional startup time.
We test the usability and flexibility of Edera by applying the design to two use cases. First, we develop a container runtime compatible with the Kubernetes specification. Kubernetes is a popular container orchestration platform for managing virtualized applications, and can be used to test Edera on large workloads. Further, Kubernetes compatibility provides interoperability with the full range of tools built by the Kubernetes community for scalability, monitoring, and more. This implementation demonstrates that Edera provides the same ease-of-use as systems with OS virtualization and can be used as a drop-in replacement for popular container runtimes.
Second, we use Edera for driver isolation, moving device drivers out of the hypervisor and into isolated zones. In particular, we develop an isolated Graphics Processing Unit (GPU) driver zone using Edera. The use of cloud GPUs is gaining popularity for the training and running of AI applications. However, bugs in GPU drivers remain common [57, 20, 54] and can be hard to detect [57, 20] leading to memory corruption in the host kernel and container escapes [11]. To reduce the risk of a GPU vulnerability impacting the kernel, we use an Edera zone to isolate GPU drivers from the hypervisor, without requiring changes to GPU drivers or hardware. This isolation prevents any misuse of the GPU from impacting the host kernel. This implementation demonstrates how Edera can be used to isolate device drivers in addition to applications. The same driver isolation model can be applied to drivers for networking, storage, or peripherals.
In summary, our contributions are:
  • We present Edera, an optimized type-1 microkernel hypervisor with strong isolation between applications.
  • We compare Edera to other virtualization systems in terms of isolation and performance, and find that it has 10.2% slower CPU speed, an average of 8.3% slower memory, and 648 milliseconds of additional startup time compared to using Docker with runc, while providing strong isolation. When compared to systems with strong isolation, Edera outperforms gVisor by all performance metrics and has comparable performance to Kata Containers on a system with virtualization extensions.
  • We implement a drop-in Kubernetes container runtime replacement that gives users the ease-of-use of modern OS virtualization systems. We further present the first system that allows for isolated AI applications on a shared GPU.
2 Background
Figure 1: A summary comparison of different virtualization designs. Note that containers in this diagram use operating system virtualization, sharing a kernel across all applications.
We provide background on the isolation and performance of virtualization techniques. A summary of virtualization designs is shown in fig. 1. We further provide background about Kubernetes and GPU isolation to motivate our use cases.
2.1 Operating system virtualization
Operating system (OS) virtualization allows multiple containers to run on a shared kernel. These containers include an application along with its runtime for independent deployment. Implementations of OS virtualization include Linux containers [36], Solaris zones [51], and Apiary [50]. Containers simplify deployment and scaling of applications by allowing multiple copies of an application to be dynamically added or removed, and a large ecosystem has been built around their use. However, there is no strong isolation between containers running on the same infrastructure.
OS virtualization uses Linux isolation techniques to prevent containers from interfering with each other, including the use of namespaces [34], cgroups [33], seccomp [35], capabilities [32], and SELinux [40]. However, as containers still operate on the same kernel, many attacks have circumvented these isolation techniques to gain access to the kernel or other containers running on the same machine [1, 24, 43, 15, 16, 9, 7, 6, 10, 8]. To prevent these container escape attacks, systems require stronger isolation.
2.2 Hypervisor virtualization
In contrast to OS virtualization, hypervisor (or virtual machine) virtualization uses a separate “guest” OS for each “virtual machine” (VM) with a hypervisor creating and monitoring multiple VMs. This provides a separate kernel for each application, ensuring stronger isolation. In a type 1 (or bare metal) hypervisor, the hypervisor runs directly on the host hardware. Examples include Xen [4] and VMware. Type 2 (or hosted) hypervisors such as VirtualBox instead run as a process on a host OS. Many hypervisors use x86 virtualization, emulating all hardware that the guest OS expects. To reduce the overhead of virtualization, hypervisors with paravirtualization [56, 4] instead require the guest OS to replace system calls with hypercalls to the hypervisor. This hypervisor-aware design avoids costly emulation of I/O hardware components. Some hypervisor virtualization systems, such as Kernel Virtual Machine (KVM), utilize virtualization extensions available on some modern Intel VT or AMD-V CPUs, allowing for more efficient hypervisor virtualization. However such virtualization extensions are not available on all CPUs.
2.3 Isolation Comparison
Unlike hypervisor virtualization, OS virtualization includes the shared kernel in the trusted computing base (TCB). By including the kernel in the TCB, OS virtualization has weaker isolation. This difference can be seen in practice in the abundance of container escape attacks on systems using OS virtualization [9, 7, 6, 10, 8, 11]. These attacks occur despite the use of Linux isolation techniques because the kernel is part of the TCB, and applications are thus vulnerable to flaws in the kernel design. The shared OS and larger TCB increases the attack surface of OS virtualization, allowing privilege escalation, side channel attacks, and information leakage between containers. Despite this, OS virtualization is common in practice because of its speed and the large ecosystem built around container orchestration technologies like Kubernetes.
2.4 Additional Isolation Techniques
Hypervisor virtualization. Several systems improve the efficiency of hypervisor virtualization. Unikernel approaches [24, 43, 19, 47, 37] use a minimized, per-application kernel to reduce the overhead of creating VMs. This approach generally (with the exception of HermiTux [47]) requires re-building images specifically for the unikernel, which can be slow and may be impossible to do for proprietary applications without source code access. A unikernel approach also means that changes to images require re-compilation of the unikernel, hampering development speed.
Hardware virtualization extensions can also be used to improve efficiency of hypervisor virtualization. Kata Containers [29], Firecracker [1], PVM [15], and RunD [22] use hardware virtualization extensions to reduce virtualization overhead. However, the reliance on virtualization extensions means these systems are not able to run on any modern CPU.
OS Virtualization. Other isolation techniques add isolation layers to OS virtualization. These include gVisor [27] which implements user-level kernel isolation, and vKernel [16] which allows for container-specific security rules. However, these systems do not achieve the isolation of hypervisor virtualization as they still contain a shared kernel.
Dedicated hardware. Other approaches to improving container isolation require dedicated hardware. Oxide computers [48] and Metalvisor [25] use dedicated hardware to more securely run containers. Constellation [26] uses a confidential VM for each Kubernetes cluster, with the cluster encrypted in an enclave during runtime.
2.5 Kubernetes
Kubernetes [31] is a system for automatically deploying, scaling, and managing containerized applications. It runs bundled applications, known as container images, in pods that run on nodes. In the OS virtualization model above, a pod is equivalent to the application and nodes are equivalent to the virtual or physical machine running the shared kernel. A container runtime manages the lifecycle of pods. Kubernetes can work with any container runtime that follows the Container Runtime Interface (CRI) specification [30]. One such container runtime is runc, a lightweight OS virtualized runtime based on Docker. The Kubernetes control plane is responsible for deploying and managing pods. This is done primarily by the Kubernetes API server, which lets users query and change the state of Kubernetes. A kubelet is an agent that runs on each node that registers the node with the API server, and ensures that all pods expected by the API server are running.
The Kubernetes ecosystem is composed of projects that add functionality, interacting with each other through defined APIs. These include container runtimes, image distribution services, and specialized control planes. The scope of this ecosystem means that any incompatible new technology will be at a huge disadvantage as it would lack the pre-built container images and services that can work with Kubernetes.
Several systems have been built to allow Kubernetes to manage VMs instead of or in addition to OS virtualized containers. This provides the usability benefits of Kubernetes to hypervisor virtualization. Many of these, including Infranetes [28], rkt [38], and virtlet [41] require virtualization extensions. RunX [39] uses a series of shell scripts to run VMs on Xen, but tooling limitations make this project incompatible with current versions of Kubernetes. These projects demonstrate the desire for hypervisor virtualization systems compatible with Kubernetes, but none meet our performance and usability requirements.
2.6 GPU Driver Isolation
GPUs are increasingly used for general-purpose computation like machine learning and high performance computing, leading to an increased demand for GPUs in multi-tenant cloud environments. Multi-tenant environments reduce the cost of using a GPU as a single GPU may be shared among many users or applications. However, as with applications running on shared infrastructure, the shared GPU presents a new attack surface that includes GPU hardware and firmware. This GPU driver firmware is a complex piece of software with many known weaknesses [57, 13] despite efforts to improve their usability and security. Amazon EC2 [2] offers GPU instances that directly pass-through access to GPUs. However, each GPU is limited to a single VM. Some systems [14, 53] use a hypervisor to allow multiple VMs to access the same GPU. In these systems, the GPU device drivers run directly on the hypervisor, increasing the TCB. To achieve stronger isolation, some systems [17, 49] propose running GPU drivers within a trusted execution environment (TEE). Similarly, Graviton [55] proposes a design for a GPU TEE which requires hardware changes to the GPU, but provides full isolation of the GPU from the OS and CPU. None of these systems ensures that compromises in the GPU are fully isolated from the hypervisor without specialized hardware or changes to GPUs.
3 Requirements
Requirement Docker Unikernel gVisor Firecracker Kata Edera Edera hwv
Eliminate shared kernel $\times$× $\times$×
Existing images $\times$×
Runs on any commodity hw $\times$× $\times$× $\times$×
Driver isolation $\times$× $\times$× $\times$× $\times$× $\times$×
Memory safety $\times$× $\times$× $\times$× $\times$×
Runtime performance $\times$× $\times$×
Startup performance $\times$× $\times$× $\times$×
Table 1: We assess existing systems against our requirements. Docker using runc is included as a baseline as it uses OS virtualization. Edera hwv shows the results for Edera running on a system with hardware virtualization extensions.
The goal of Edera is to create strong container isolation, while ensuring usability. Usability means both the ability to take advantage of the large ecosystem built around containers, and the ability to adopt Edera without affecting runtime performance. To achieve this, our requirements are as follows.
Eliminate the shared kernel. The system must support secure multi-tenancy, with the container acting as a security boundary. This means removing the kernel from the TCB so that a bug in the kernel is not sufficient to escape the isolation.
Use existing images. Existing container images must run without modification. Some images have proprietary code which cannot be accessed by users to re-build the image. Even when the source is available, re-building images is time consuming.
Run on any commodity hardware. The system must be able to run on common, commercially available hardware without relying on specialized hardware like virtualization extensions or TEEs. Virtualization extensions are not available on all CPUs or on all commercial cloud infrastructure, and can be much more expensive when available. For example only 7% of AWS instance types include virtualization extensions.
Driver isolation. The system should be able to isolate hardware drivers, including those for GPUs, from the host kernel. In other words, a driver compromise should not be able to impact the hypervisor or host machine.
Memory Safety. The system should be written in a memory safe language to reduce the attack surface of the hypervisor. Memory safety vastly reduces the instance of memory leaks and memory corruption attacks.
Runtime performance. The system’s runtime performance must be within 15% of running runc on Docker. This ensures that the performance is similar to that of systems using OS virtualization.
Startup performance. The system’s startup performance must be within 10% of that of runc on Docker. While the system startup occurs less often than runtime (just once per application), it is still important to have a fast startup time for provisioning new containers and scaling the number of copies up and down.
We analyze several systems discussed in section 2, and show how they compare to these requirements in table 1.
3.1 Threat model
In multi-tenant environments, arbitrary workloads from third parties are run on shared infrastructure. The third party may be a company running their infrastructure for several customers, or a customer running their own containers. The infrastructure allows customers to run their own code (often in the form of Kubernetes containers), but the operator does not want customers to gain access to the underlying infrastructure or other customer workloads. Given this environment, for our threat model we assume an attacker can:
  • Run an application in a zone on a multi-tenant machine.
  • Have root access to the Linux kernel in that zone.
  • Run any Linux system call in the zone.
  • Run commands on a shared GPU from the zone.
In this environment, an attacker succeeds if they are able to:
  • Access another zone, the hypervisor, or root zone.
  • View memory, CPU usage, or other information about containers running in other zones.
  • Gain access to a running zone that they do not control.
The following are out of scope:
  • Supply chain attacks on the hypervisor or root zone.
  • Side channel attacks, for instance those that use the temperature of the host machine. These attacks are unavoidable on shared infrastructure, and are present in any virtual machine environment.
4 Design
Edera is a type-1 paravirtualized hypervisor that runs workloads in zones, which are fully functional VMs that run standard Linux for container workloads. Edera has a root hardened zone that manages the hypervisor, passing along memory and CPU from the host machine. This root hardened zone can monitor for anomalies and shut down other zones if needed. The microkernel hypervisor controls the memory, page table, and CPU usage of zones. Each zone is scheduled like a user-space process within the hypervisor. Figure 2 provides an overview of Edera.
Figure 2: An overview of how Edera isolates zones.
4.1 Hypervisor
The Edera hypervisor is based on Xen’s hypervisor design [4], with a few key differences and additional features. For brevity this section focuses on the novel features of the Edera hypervisor. The Edera hypervisor is a microkernel hypervisor that acts as the host kernel and can run directly on the host machine. By using a microkernel, Edera minimizes the attack surface of the hypervisor. Research has shown that microkernels have fewer bugs, as many bugs come from less used pieces of the kernel [21].
Edera works in paravirtualization (PV) mode, with the hypervisor managing the CPU, MMU hardware page table, and network. There is also support for paravirualization with hardware virtualization (PVH) mode, which takes advantage of virtualization hardware for performance improvements. However, as PVH mode requires virtualization hardware, examples in this paper use PV mode unless otherwise stated. The page table mapping used by Edera is the same as in Xen, with the hypervisor giving guest OSs read-only access to virtual page tables that the hypervisor maps to the MMU hardware page table. CPU scheduling in the Edera hypervisor is more dynamic than the scheduling supported by Xen, with the resource manager able to dynamically pin and unpin CPUs for each zone. This dynamic configuration allows for more efficient resource allocation on the host machine as zones can be assigned CPUs and memory pages as needed, rather than these resources remaining static after zone creation.
4.2 Guest kernel
Each zone launched by the hypervisor runs a guest OS kernel. Using PV, the hypervisor runs in ring 0, and all guest OSs run in a less privileged ring (ring 1 for x86 and ring 3 on x86-64) [44]. As most OSs expect to run in ring 0, the guest OS must be aware that it is running in a virtualized environment and support hypercalls. These hypercalls [4] are used for privileged operations that need to run in ring 0 and pass along those privileged system calls to the hypervisor. This is supported by many OSs through a PV mode.
Edera supports any Guest OS that has a PV mode, including most Linux distributions. For ease of use with Kubernetes, we create a custom guest OS kernel that is distributed as an OCI image. Thus, a user can specify this image in their Kubernetes configuration and automatically run their existing containers.
4.3 Inter-domain messaging
Edera uses inter-domain messaging (IDM) to communicate with processes running in zones. IDM uses a Xen byte channel [42] to send protobuf messages to other zones using the transport layer. Protobuf (shortened from protocol buffers) is a platform-neutral mechanism for serializing structured data. Its design includes protection against overloading the pipe by using bound checks and other limits to ensure safe operation.
IDM is used to send commands or processes to the zone to be run (such as through the Kubernetes utility kubectl exec),in addition to sending logs and metadata back to the hypervisor. Any logs received from the IDM are considered untrusted as they come from a potentially-compromised user zone. However, they can be used to detect misbehavior in the zone.
4.4 Monitoring
An init process in every zone starts when the zone is provisioned and runs a background process that monitors the VM. The init process sends all monitored events over the IDM to be logged and collected. If the init process is disabled, the hypervisor would mark the zone as not responding and deprovision it. As the init process runs on the zone, any compromise of the init process itself would only impact that zone.
Edera’s monitoring happens outside of the zone and uses information from init, CPU and memory usage, and network traffic passed through the hypervisor. As guest kernels do not have access to ring 0 without going through the hypervisor, all privileged operations can be limited and monitored. All network packets leaving the VM are monitored when they are passed through to the hypervisor, so a VM cannot send malicious traffic that bypasses the monitoring. This monitoring can be used to manage resource usage as well as to perform detection and response for processes running on the VM. For example, a misbehaving zone could be quarantined and monitored to understand malicious activity.
5 Implementation
Figure 3: An overview of the services used by Edera.
We implemented Edera in Rust and MISRA C using the design from section 4 for both Kubernetes compatibility and GPU isolation. By implementing it in Rust and MISRA C, we avoid many memory safety bugs while achieving high performance. The implementation includes the hypervisor, root hardened zone, and several services that allow use with Kubernetes for cluster orchestration. These services include a Container Runtime Interface (CRI) compatible with the OCI runtime specification, daemons for zone monitoring, and a kube-apiserver for managing pods.
To start, we create a new set of Xen tooling in Rust to replace libxl. This tooling simplifies starting a new Xen domain by reducing the number of configuration files and introducing a Rust API to replace the command line tools used by libxl. On top of this new Xen tooling, we introduce several services.
5.1 Kubernetes
We will walk through the system startup and running of a single Kubernetes pod in an Edera zone to illustrate the services used by the implementation, shown in fig. 3. First, the machine starts the microkernel when it is booted. This microkernel starts the hypervisor, which also acts as a Kubernetes host. The hypervisor launches protect-orchestrator, a management component that ensures that all other daemons are running. These daemons are protect-network, protect-daemon, and protect-cri.
protect-daemon acts as the virtual machine monitor for the zones, keeping track of which zones are running each workload and managing drivers for the zones, while leaving resource management to the hypervisor. This is the daemon that creates, modifies, and removes zones. To do so, it manages an embedded key value store that tracks what is running in each zone. In this store each zone is identified by a uuid that is assigned on zone creation. In addition, protect-daemon manages the drivers for each zone, including the IDM. Using IDM, protect-daemon can pass commands from the API server (though protect-cri) to the init process in the zone. It can also receive messages from the zone, and pass these along to protect-cri.
protect-orchestrator performs the initial system setup, and maintains running processes. It provides updates and health checks for running daemons including protect-daemon, protect-network, and protect-cri. In addition, it provides observability for all services running in the hypervisor, through tools like Prometheus [3].
protect-network performs all packet routing and monitors network activity. Zones have a basic ethernet interface, but cannot perform packet routing without going through protect-network. protect-network receives tasks from zones through protect-daemon, then performs them on behalf of the zone. It writes packets to the networking layer and can dial sockets for zones. Zones do not have direct network access, so no packet gets routed without being seen by protect-network. This ensures that Edera can monitor all network activity.
protect-cri provides a layer on top of the Kubernetes CRI, which is the protocol for communication between a kubelet and the container runtime. It can be used with any OCI-compliant CRI, and converts requests from the CRI to daemon requests for protect-daemon. protect-cri reconciles the state between protect-daemon and the CRI to ensure that the zones running on protect-daemon match those expected by the CRI. It also transfers messages from zones through protect-daemon to other Kubernetes processes. protect-cri supports running some Kubernetes nodes in Edera protect zones and others on different infrastructure. To do so, the protect-cri coordinates with containerd to run the non-Edera nodes.
Once the system is booted, a user may launch a pod from protect-cri. This triggers the Kubernetes kubelet [18], which registers the pod with the kube-apiserver and launches a zone.
Summary. In summary, Edera provides several services: protect-daemon, protect-orchestrator, protect-network, and protect-cri to manage zones and facilitate communication with Kubernetes. By routing all network traffic through protect-network and all IDM communication through protect-daemon, we ensure that all traffic can be monitored.
Figure 4: An overview of the use of Edera for GPU isolation. GPU drivers are moved from the hypervisor into a zone so that any misbehavior cannot effect the hypervisor or other zones.
5.2 Driver isolation
In addition to using Edera to interoperate with Kubernetes, we build device driver isolation using Edera zones. We demonstrate that Edera can be used for strong driver isolation that prevents device driver bugs from impacting the hypervisor. This driver isolation can be used along with the Kubernetes compatibility described above.
To illustrate driver isolation, we discuss the isolation of GPU drivers. Usually, drivers for GPUs and other hardware devices are run on the host kernel directly, then a sub-driver is exposed to each VM. If a VM requests an operation of the device, this operation is passed through to the driver on the host kernel. However, with this design, any bug in the device driver can lead to a corruption of the host kernel. Device drivers can be many gigabytes of proprietary code and have had bugs discovered in the past, with more than 20 CVEs in 2024 [45]. Further, tight integration between GPUs and CPUs provide a large attack surface for GPU bugs [57].
Instead, Edera runs device drivers in a zone with the GPU attached to this zone, as shown in fig. 4. Applications run in their own zone and interact with this zone through an NVIDIA vGPU [46] that communicates with the GPU driver zone. The driver zone can be configured to passthrough the entire GPU to a single zone, or to share a GPU among multiple zones using partitioning. If any GPU user exploits a driver bug, only this GPU zone, and not the hypervisor or other user zones, will be compromised. Further, the zone can be configured to wipe GPU memory between different users of the GPU to prevent data leakage.
This same technique for driver isolation can be applied to other drivers. For example, network drivers managed by protect-network in fig. 3 can be moved into an Edera zone. Thus, network operations would be sent to this zone instead of the hypervisor. This moves these drivers out of the TCB and ensures that any vulnerability in the driver code can only be exploited within the zone and will not effect the hypervisor.
6 Analysis
In order to analyze Edera, we compare it to other container isolation techniques in terms of performance and security.
6.1 Performance
We compare the performance of Edera under both to several existing systems. Docker using runc acts as a baseline metric for the performance of operating system virtualization. gVisor and Kata Containers (Kata) allow a comparison with other systems with strong isolation. We especially compare Edera in PV mode to Docker and gVisor, as none of these systems require virtualization extensions, while comparing Edera PVH with Kata to compare performance when such extensions are available. We thus run all systems on an OVH virtual machine with virtualization extensions, and run Docker and gVisor benchmarks both on bare metal and on the Edera hypervisor to separate out the performance impact of the Edera hypervisor from that of running benchmarks in Edera zones.
To assess performance, we evaluate both startup and runtime. Within runtime performance, we look at the CPU and memory usage, as well as the speed of various system calls. To estimate real-world performance, we also evaluate the runtime of each system for some example workloads. We run all tests 5 times and present the average of these runs, with the standard error shown in all graphs.
Figure 5: A comparison of the memory speed in MB/s using metrics from both sysbench and perf-bench. Higher is better.
Figure 6: A comparison of the CPU speed in events per second using metrics from sysbench. Higher is better.
Startup time. We first examine the startup time of each system, shown in table 2. Edera takes 0.77 seconds to startup in PV mode, compared to 0.97 seconds in PVH mode. This is a slower startup time than Docker and gVisor, which start up in 0.18 and 0.2 seconds respectively, but faster than Kata Containers, which starts up in 1.93 seconds. However, Edera zones can be started before they are needed and kept running in an idle state until use. In this way, the user does not need to wait for the 0.77 second startup time each time they use a zone. These idle warm zones can initially be allocated no memory or CPU resources, then resources can be dynamically allocated once the zone becomes active due to the flexibility of the Edera hypervisor.
system startup time (ms)
Docker 177.4
Docker (bare metal) 203.8
gVisor 281.8
Edera-PV 765.8
Kata Containers 1934.2
Edera-PVH 968.6
Table 2: The startup time for each system.
Memory and CPU benchmarks. Figure 5 shows the memory usage from sysbench, a benchmarking tool that performs reads and writes to a data buffer, and the speed of the memset and memcpy operations from the perf-bench utility. Edera outperforms the other systems on both sysbench tests, although it is only slightly faster than Docker and Kata on bare metal (.03% and 1.8% respectively). Edera also equals or outperforms Docker, gVisor, and Kata for both perf-bench operations. Note that the hypervisor does not make a large difference for memory operations, with Docker performing similarly both on bare metal and on the Edera hypervisor.
Figure 6 illustrates the performance of the sysbench CPU benchmark. For this test sysbench monitors CPU usage while checking prime numbers up to 150,000 over 4 threads. Edera is slightly slower than both Docker and Kata on bare metal (.9% and .7% slower respectively). Docker runs much slower on the Edera hypervisor than on bare metal, illustrating that some of the CPU performance lost through the use of vCPUs in the Edera hypervisor is made up through the CPU management by Edera zones. Edera benefits from using paravirtualization for resource management, as it allows for faster memory mapping and CPU scheduling by using a software interface instead of full hardware emulation. Edera has a full view of the hardware, allowing smart optimizations like co-locating pages in a zone.
Figure 7: A comparison of the operations per second of system calls. Note that this graph is shown on a logarithmic scale. Higher is better.
Figure 8: A comparison of the system call latency. Note that this graph is shown on a logarithmic scale. Lower is better.
Figure 9: A comparison of runtime of kcbench in milliseconds. Lower is better.
Figure 10: A comparison of runtime of nginx in milliseconds. Note that this graph is shown on a logarithmic scale. Lower is better.
System calls. Next, we evaluate the speed of executing several popular system calls. The mechanism by which system calls are passed to the host system differs widely between the systems we evaluate, so we determine how these approaches impact performance. Figure 7 shows the operations per second of each system running the fork(), execve(), getpgid(), and epoll() system calls based on results from perf bench. This graph shows that Edera outperforms both Docker and gVisor for the execve and epoll operations, and outperforms gVisor for fork and getpgid. Edera avoids much of the indirection used by gVisor, leading to much faster performance. Kata outperforms Edera for fork(), execve(), and getpgid(). However, in PVH mode, Edera performs fork() and execve() similarly to Kata. Both Kata and Edera PVH use virtualization extensions to achieve this higher system call performance. Figure 8 further shows the latency of fork(), execve(), and getpgid(). The latency of Edera in PVH mode is similar to that of Docker and Kata. Docker also has lower latency when run under the Edera hypervisor, illustrating the scheduling performance of the hypervisor.
Real-world workloads. Finally, we evaluate the runtime of workloads in each system to understand how they perform in real-world environments. These runtimes are shown in fig. 9 and fig. 10. We first discuss the time to build a kernel in each system using kcbench. Edera’s kernel build time is 5% slower than Docker, 59% faster than gVisor, and 52% faster than Kata. This kernel build benchmark represents a more realistic workload with several different system calls and operations. Next, we evaluate the runtime of a simple nginx server using Apache Bench. We find that Edera outperforms gVisor by a factor of 10 (18 ms compared to 220 ms for Docker), and performs comparably to Docker (15 ms). Edera-PVH has slightly faster performance than Kata according to this metric, although Edera-PV is slightly slower. The use of a webserver like nginx is a common application for cloud applications.
6.1.1 Summary
In summary, Edera outperforms gVisor in all performance metrics tested with the exception of startup time, while overall performing only slightly worse than Docker and similarly to Kata. This means that Edera achieves strong application isolation with minimal impact on performance even for systems that do not have access to virtualization extensions. The system startup time performs the worst, with 0.77 seconds of startup time for Edera compared to 0.2 seconds for Docker. When virtualization extensions are available, Edera in PVH mode has improvements to system call runtime over PV mode, and outperforms Kata on memory and CPU operations.
6.2 Security
Attack Mitigation Protected?
CVE-2022-0185 No shared kernel
CVE-2022-0492 No shared kernel
CVE-2022-0811 (cr8escape) No shared kernel
CVE-2022-0847 (Dirty Pipe) No shared kernel
CVE-2022-23648 No shared kernel
CVE-2024-0132 Driver isolation
CVE-2024-21626 (Leaky Vessels) Runtime exploits can’t reach host
Table 3: Overview of recent container escape CVEs and how they are mitigated by Edera.
Next, we evaluate the security of Edera using our requirements from section 3 and previous container escape attacks.
6.2.1 Requirements
Edera protect meets the requirements from section 3:
  • Eliminate shared kernel. Edera uses a type-1 hypervisor to isolate containers into zones, with each zone containing its own guest operating system. This separation eliminates the shared kernel used by OS virtualization.
  • Use existing images. Edera is compatible with existing OCI images, and protect-cri complies with the Kubernetes CRI specification for seamless use by existing systems. This reduces the barriers to use by ensuring that existing applications, even those using proprietary software, can be run in Edera without modification.
  • No specialized hardware required. Edera works on all commercially available hardware, even that without virtualization extensions. It does so by using paravirtualization on guest kernels to allow full use of Linux system calls without the need for hardware virtualization.
  • Driver isolation. Edera allows hardware drivers to run in isolated zones, protecting the hypervisor in the event of a driver compromise.
  • Runtime performance. Edera achieves faster runtime performance than gVisor by all metrics and performs within 15% by all metrics when compared to Docker with the exception of getgpid. This meets the requirement specified in section 3. When run in PVH mode, Edera achieves even better system call and workload performance.
  • Startup performance. Edera adds 0.56 seconds of startup time compared to our baseline of Docker with runc. Although this startup time does not meet the requirement from section 3, it can be further reduced by using warm zones to meet this requirement.
6.2.2 Preventing container escapes
Next, we evaluate Edera’s security by examining how it can be used to prevent container escapes. We evaluate seven container escapes from 2022-2024 to determine how Edera fares against these attacks. We focus on vulnerabilities in Linux in this timeframe, as these could impact both Edera and other isolation techniques and leave out analysis of vulnerabilities in other isolation techniques. Table 3 shows the results.
Findings. We find that Edera prevents all of these attacks through a combination of having no shared kernel and driver isolation. The prevention mechanism is labeled in table 3 for each attack. Without a shared kernel, attacks on capabilities, namespaces, and other OS isolation techniques do not give an attacker access to the host machine. This means that any kernel exploit that previously could cause a container escape will only impact the guest OS on Edera. As such, Edera is not vulnerable to any of these past container escape attacks and is further not vulnerable to this class of attack in the future. We provide in-depth explanations of three of these CVEs to illustrate how Edera prevents them.
CVE-2022-0492. This vulnerability allows a user of cgroup release_agent to bypass cgroup’s namespace isolation, thus escalating privileges and causing a container escape. Edera does not use cgroup for isolation, instead placing the entire guest kernel in an Edera zone. As such, this vulnerability in cgroup cannot be used to escape an Edera zone.
CVE-2022-0847. This vulnerability, known as “Dirty Pipe,” was caused by the improper initialization of pipes. Pipes could contain stale values, allowing an unprivileged user to write to read-only pages in the page cache and escalate their privileges. Edera provides guest OSs with read-only access to virtual page tables, which the hypervisor then maps to MMU hardware pages. As each zone only has access to its own virtual page table, it a user in a zone is unable to write to pages in the hypervisor or other zones.
CVE-2024-21626. A vulnerability in runc, known as “Leaky Vessels,” uses a leak of an internal file descriptor to create a container with a working directory in the host filesystem namespace. This can be used to gain access to the host filesystem, or create a privilege escalation by re-writing files on the host filesystem. In Edera, guest OSs do not share a filesystem with the hypervisor, so even if a guest OS had access to a hypervisor file descriptor, they would not be able to access or overwrite files on the hypervisor.
Summary. These vulnerabilities illustrate the security gains of hypervisor isolation. By eliminating the shared kernel, including shared access to pages and the filesystem, Edera is able to prevent container escapes. Edera can isolate vulnerabilities in the kernel or drivers to limit their impact, ensuring that only resources within the zone can be affected.
7 Discussion
We discuss addition considerations when using Edera, and opportunities for future work.
7.1 Zone Flexibility
Edera allows for flexibility through dynamic resource allocation. Memory and CPU allocations for a zone can be dynamically updated, allowing for more efficient use of these resources and greater flexibility for changing zone requirements. If a zone starts with 8GB of memory but is running out, the zone can be given additional memory without interfering with the workload. In most existing systems, allocating more memory requires re-creation of the container. Further, a Kubernetes container running in a zone can be dynamically moved to another machine without interrupting the workload. This is possible because the zone is independent of the host machine and so can be moved to other hardware if needed.
7.2 Usability
The good performance of Edera will allow organizations to use security mechanisms for all containers, not just those related to security. For performance reasons, many organizations use OS virtualization for most operations, and a system with stronger isolation (like gVisor or Kata Container) for certain, security critical applications. This segmentation of operations leaves most workloads vulnerable to container escapes. Edera obviates the need for this segmentation, providing strong security for all applications.
Edera achieves this by providing a drop-in replacement for the Kubernetes runtime so that it can be used for isolation without sacrificing usability at all other layers of Kubernetes. To demonstrate, consider a simple Kubernetes configuration:
apiVersion: v1
kind: Pod
metadata:
  name: leaky-vessel
  namespace: test
spec:
  containers:
  - name: leaky-vessel
    image: example/leaky-vessel:0.1
    imagePullPolicy: Always
    env:
    - name: SUPER_ORCHESTRATOR_SECRET
      value: "this-is-fine-leaky"
By adding a 2 line annotation to this configuration, the application is now protected by Edera (duplicate configuration is removed for brevity):
metadata:
  annotations:
    dev.edera/kernel: \
    "ghcr.io/edera-dev/linux-kernel:latest"
    dev.edera/memory: "600"
spec:
  runtimeClassName: edera
7.3 Future work
Opportunities for future work include:
  • Further optimizations to improve Edera’s startup and runtime performance.
  • Expanding Edera’s driver isolation to non-NVDIA GPUs. Other GPUs have different supported drivers, and do not support vGPU used by our implementation.
  • Hardening the Edera hypervisor through the use of hardware enclaves or support for confidential computing. For example the startup process could be verified by hardware keys stored in an enclave.
  • While Edera limits kernel exploits to a single zone, it could be expanded to provide protection within the zone through the use of monitoring. A protocol like expanded Berkeley Packet Filtering (eBPF) could be used in the hypervisor to detect anomalies and prevent exploits running in a zone. Further, monitoring of known compromised kernels could be used for compromise detection.
8 Conclusion
We present and evaluate Edera, a type-1 hypervisor that has performance comparable to popular operating system virtualization technologies while providing strong isolation. We implement Edera for both a Kubernetes compatible container runtime, and for driver isolation to demonstrate how it can work within existing infrastructure. We find that Edera outperforms gVisor in CPU performance, memory performance, and under real-world workloads, while only adding 0.48 seconds of startup time without the use of virtualization extensions. It further achieves comparable performance to Docker and Kata Containers. By eliminating the shared kernel, Edera is able to prevent all tested container escape attacks.